# EA-KD: Entropy-based Adaptive Knowledge Distillation

### v2

@@ 4.1 Experimental Setup
Datasets. We used CIFAR-100 [12] (50k training, 10k validation images; 100 classes) and Tiny-ImageNet [14] (100k training, 5k validation images; 200 classes) to evaluate our method for image classification, and used MS-COCO [18] (118k training, 5k validation images; 80 classes) for object detection. Furthermore, for LLM distillation, we employed five instruction-following datasets following [7]: Dolly [4], SInst [34], Vicuna [3], S-NI [35], and UnNI [10].
Table 3: Results on Tiny-ImageNet. Top-1 (first row) and top-5 accuracy (second row) for both EA-methods and their original versions using ResNet32$\times$4 teacher and ResNet8$\times$4 student are reported.
Teacher Student KD EA-KD MLD EA-MLD MLD+LS EA-MLD+LS FCFD EA-FCFD Avg. $\Delta$
64.41 55.25 56.00 59.39 +3.39 61.91 62.65 +0.74 56.45 61.93 +5.48 60.12 60.51 +0.39 +2.50
85.07 79.62 79.64 82.40 +2.76 83.77 84.47 +0.70 81.15 84.35 +3.20 82.90 83.22 +0.32 +1.74
Table 4: Results on Tiny-ImageNet with Transformers. Accuracy (%) for EA-KD and KD using transformer-based teachers and a ResNet8$\times$4 student. Each result is averaged over three runs.
@@
Table 5: Results on MS-COCO. Performance of EA-KD and KD using Faster-RCNN with FPN. Metrics include AP (overall), AP50, and AP75.
@@
Implementation Details. The EA-KD and other variants were evaluated on various CNNs (VGG [29], ResNet [8], WideResNet [36], MobileNet [27], ShuffleNet [23]), vision transformer teachers (ViT [6], DeiT [32], Swin [21]), and GPT2 [25] for LLM distillation. We used the training settings of [31] for vanilla KD and the respective settings of each KD framework for CIFAR-100 and Tiny-ImageNet, [1] for MS-COCO, and [37] for LLMs. We set $T^{\prime}=3$ based on our ablation study, except $T^{\prime}=2$ for transformer teachers. All results averaged over five runs, and analyses were conducted using a ResNet32$\times$4 teacher and a ResNet8$\times$4 student on CIFAR-100.
@@ 4.2 Results
Tiny-ImageNet. Tab. 3 presents results on Tiny-ImageNet across SOTA frameworks enhanced with EA. For instance, EA-KD significantly improves top-1 accuracy by 3.39% and top-5 accuracy by 2.76% over standard KD, achieving performance close to FCFD with less than a 1% difference. These results highlight EA’s effectiveness on a larger and more class-diverse dataset. Further, Tab. 4 highlights EA-KD’s ability to outperform KD when distilling from transformer-based teachers to a CNN student. While standard KD stopped improving the student when Swin-B was used as the teacher, EA-KD continued to enhance student performance, mitigating the capacity gap by allowing better teachers to produce better students.
@@
Table 6: LLM Distillation Results. Rouge-L [17] scores averaged over five seeds for KD, EA-KD, RKLD, and JSD on each dataset. SFT denotes a student fine-tuned on the dataset.
@@ 4.3 Empirical Analysis
(a) KLD for Top and Bottom 10% Entropy Samples
(b) Raw KLD and Training Accuracy (%)
Figure 5: Comparison of KD and EA-KD During Training. EA-KD enhances focus on high entropy samples, leading to a more effective and efficient learning process, as reflected in the overall raw KLD and training accuracy.
Ablation Study. Two experiment were performed to evaluate the sensitivity of $T^{\prime}$ and the effect of each reweighting components in $w_{\text{EA}}$. (i) Tab. 7 shows that a $T^{\prime}$ of 3 consistently delivers optimal performance across model combinations, highlighting its robustness in reflecting sample value during distillation. (ii) As shown in Tab. 8, the single-sided factor $w_{\text{base}}$ notably boosts baseline performance by emphasizing valuable samples, while $w_{\text{interact}}$ integrates teacher-student dynamics but may overly rely on the student responses, leading to limited gains. By combining these strengths, $w_{\text{EA}}$ achieves superior improvement by prioritizing valuable samples and adapting to the student’s learning dynamics, enabling a precise and tailored focus.
@@
KD [9] 73.33 $\uparrow$ 75.14 $\uparrow$ $\uparrow$ 74.76 $\uparrow$ $\uparrow$ 75.27 $\uparrow$
@@
FCFD [19] 76.62 $\uparrow$ 77.37 $\uparrow$ $\uparrow$ 77.42 $\uparrow$ $\uparrow$ 77.44 $\uparrow$
Refined Knowledge Focus and Alignment. Fig. 5(a) illustrates the KLD for the top and bottom 10% entropy samples throughout training for both KD and EA-KD, with Fig. 5(b) depicting the overall raw KLD and training accuracy. Notably, vanilla KD’s uniform distillation causes low-entropy samples to dominate the distillation loss, failing to transfer the more critical knowledge from higher-entropy samples, as discussed in Sec. 3. Fig. 2(a) further illustrates this, where KD shows increasingly larger accuracy gaps with the teacher on higher entropy samples. In contrast, EA-KD refined the student’s focus on high-entropy samples and efficiently reduced the raw KLD (Fig. 5), achieving improved mimicry and generalization with more consistent alignment to the teacher across entropy segments (Fig. 2(a)).
Student’s Unique Perspective. Fig. 6 compares the t-SNE visualizations [33] of KD and EA-KD with the high $H^{\mathcal{S}}$ and $w_{\text{EA}}$ samples highlighted. Similar to Fig. 2(b), high $H^{\mathcal{S}}$ samples also lie near decision boundaries in the KD-student. This suggests that, despite having a different view from the teacher (Fig. 3), the student’s evolving capacity enables $H^{\mathcal{S}}$ to capture the samples crucial for its own learning progress [^3] See Sec. A.1 for further t-SNE visualizations showing the student’s evolving sample focus over training epochs.. In EA-KD, integrating both $H^{\mathcal{T}}$ and $H^{\mathcal{S}}$ in $w_{\text{EA}}$ allows us to harness the complementary strengths of the teacher’s informed assessment and the student’s dynamic learning needs. Consequently, EA-KD achieves superior class separability as indicated by a higher Calinski-Harabasz (CH) index.
@@
Figure 6: t-SNE of Students. Top 10% high $H^{\mathcal{S}}$ samples for KD (left) and $w_{\text{EA}}$ samples for EA-KD (right) are highlighted with triangles. High $H^{\mathcal{S}}$ samples cluster near decision boundaries, underscoring their learning value for the student. EA-KD achieves better class separability as indicated by a higher CH index.
@@
The enhanced generalization in EA-DKD also extends to the robustness of its hyperparameter. DKD was introduced to decouple target class KD (TCKD) and non-target class KD (NCKD) in vanilla KD, balancing them with hyperparameters $\alpha$ and $\beta$ [38]. While $\alpha$ remains stable around 1.0, adjusting $\beta$ from 1.0 to 10.0 leads to fluctuating performance [38], with a variance of 0.31 (see Fig. 8). Notably, EA-DKD effectively reduces this fluctuation to a variance of 0.10 and consistently improves the performance. Overall, the enhanced generalizability and robustness may be attributed to the synergy between DKD and EA in handling different aspects of knowledge transfer: DKD prevents NCKD from being overshadowed by TCKD at the class level, while EA ensures valuable samples—often rich in NCKD—are not dominated by simpler ones at the sample level. Additionally, the dynamic focus in EA compensates for the static nature of $\beta$. Together, EA-DKD facilitates a more balanced and efficient knowledge transfer process, both class-wise and sample-wise.
Computational Efficiency. Fig. 1 illustrates the efficiency of EA-methods, showing significant performance improvements on various SOTA KD frameworks with negligible cost. This remarkable computational efficiency, combined with its streamlined integration into both logit- and feature-based KD frameworks, suggests the potential of our method as a versatile and accessible enhancement for KD.
@@ 5 Conclusion
In this paper, we examined existing KD methods from a novel perspective, revealing the limitation of their inherent uniform distillation scheme, which often hinders the transfer of valuable, high-entropy samples that contain critical knowledge. To overcome this, we introduced EA-KD, a plug-and-play KD approach that dynamically reweights the distillation loss, directing the learning process toward valuable samples. EA-KD consistently improves SOTA KD methods across image classification, object detection, and LLM distillation, all with negligible cost. We believe EA-KD showcases a great paradigm of the meticulous handling of knowledge transfer, offering sample-wise adaptability while integrating the student’s evolving learning dynamics into the distillation process.
@@
Appendix A Additional Analysis and Visualizations
This document provides additional analysis to complement the main paper. It includes a detailed examination of high student entropy samples, loss landscape comparisons for KD and EA-KD, and t-SNE visualizations for various KD frameworks and their EA-enhanced variants.
A.1 Analysis of High Student Entropy Samples
Complementary to the teacher entropy analysis in Sec. 1, this section emphasizes the critical role of high student entropy samples in the adaptive reweighting process of EA-KD. As shown in Fig. 1(a), high student entropy samples correlate with larger teacher-student accuracy gaps similar to the teacher entropy segments (Fig. 2(a)). Additionally, Fig. 1(b) presents the students’ t-SNE visualizations across training epochs for KD and EA-KD, with high-entropy samples ($H^{\mathcal{S}}$) and reweighted samples ($w_{\text{EA}}$) highlighted, respectively. Notably, these high-entropy samples also cluster near decision boundaries similar to the teacher’s t-SNE (Fig. 2(b)). Unlike the teacher, however, the student’s top-entropy samples shift dynamically over training. For instance, in KD, the yellow class at epoch 40 is located near the center of most clusters and contains numerous high-entropy samples. By epoch 120, high-entropy samples are more prevalent in the green and purple classes as they shift closer to the center. With this dynamic integrated into $w_{\text{EA}}$, EA-KD continuously adapts to the student’s evolving learning needs throughout training, complementing the static nature of $w_{\text{base}}$ and fostering a tailored knowledge transfer.
@@ A.1 Analysis of High Student Entropy Samples
Figure A1: High Student Entropy Analysis in KD and EA-KD. (a) Higher student entropy samples correlate with larger accuracy gaps in KD. EA-KD demonstrates closer alignment with the teacher, particularly in high-entropy regions. (b) High $H^{\mathcal{S}}$ samples consistently cluster near decision boundaries, reflecting the student’s real-time learning needs. EA-KD adapts to this dynamic, achieving enhanced class separability over epochs.
A.2 Loss Landscape Analysis for KD and EA-KD
We further analyze the robustness of KD and EA-KD by examining their loss landscapes across training epochs. As shown in Fig. A2, EA-KD achieves a consistently smoother loss surface compared to KD (e.g. $\sigma^{2}$ of 217.68 vs. 305.21 at epoch 240) and larger Area@1.6 values. Notably, KD’s Area@1.6 metric remains 0 at epochs 40 and 120, indicating the absence of low-loss regions. In contrast, EA-KD achieves values of 0.01 and 0.05, respectively, underscoring its improved learning efficiency by enabling the student to assimilate the key knowledge. These results highlight EA-KD’s ability to ensure a more efficient and effective optimization, facilitating enhanced generalizability compared to standard KD.
A.3 t-SNE of KD Frameworks and EA-methods
The t-SNE visualizations of students from various KD frameworks and their EA-method variants, alongside the teacher’s representation, are shown in Fig. A3. EA-methods consistently exhibit more distinct and well-defined class clusters compared to their respective baselines, as evidenced by higher CH indices. Remarkably, the SOTA EA-MLD+LS achieves the highest CH index of 814.96, indicating its superior performance and closer alignment with the teacher. Furthermore, although EA-DKD ranks fourth in performance (Tab. 2), it demonstrates the second-best class separability, underscoring its enhanced robustness as discussed in Sec. 4.3. These findings highlight EA-KD’s versatility in enhancing class separability and improving knowledge transfer across diverse KD frameworks.
@@ A.2 Loss Landscape Analysis for KD and EA-KD
Figure A2: Loss Surface and Differences in Area@1.6 for KD and EA-KD Students Across Epochs. Contour plots illustrate the loss landscapes of KD (first row) and EA-KD (second row) across training epochs. The line plot (lower right) tracks the differences in Area@1.6 between the teacher and students over epochs. EA-KD exhibits a more stable and robust loss surface, with greater Area@1.6 earlier in training, signifying a more efficient learning process.

### v3

@@ 4.1 Experimental Setup
Datasets. We used CIFAR-100 [16] (50k training, 10k validation images; 100 classes), Tiny-ImageNet [18] (100k training, 5k validation images; 200 classes), and ImageNet [6] (1.28M training, 50k validation images; 1,000 classes) to evaluate our method for image classification. For object detection, we adopted MS-COCO [25] (118k training, 5k validation images; 80 classes). Furthermore, for LLM distillation, we employed five instruction-following datasets following [9]: Dolly [5], SInst [42], Vicuna [4], S-NI [43], and UnNI [12].
[…]
Teacher Student KD EA-KD MLD EA-MLD MLD+LS EA-MLD+LS FCFD EA-FCFD
64.41 55.25 56.00 59.39 +3.39 61.91 62.65 +0.74 61.36 62.41 +1.05 60.12 60.51 +0.39
Table 4: Results on ImageNet. Accuracy (%) of the ResNet34 teacher and ResNet18 student, averaged over three runs.
Teacher Student KD [11] EA-KD KD+LS [38] DKD [47] EA-DKD DKD+LS [38] EA-DKD+LS PAD [46]
73.31 69.75 71.03 71.79 +0.76 71.42 71.70 71.96 +0.26 71.88 71.99 +0.11 71.71
[…]
@@
Table 6: Results on MS-COCO. AP (overall), AP50, and AP75 are reported using Faster R-CNN with FPN.
@@
[…]
@@ 4.2 Results
Tiny-ImageNet and ImageNet. Tab. 3 and Tab. 4 show that EA-methods consistently outperform their baselines on both Tiny-Imagenet and ImageNet. On Tiny-ImageNet, EA-KD improves KD by 3.39% and achieves performance comparable to the novel FCFD [26]. On ImageNet, EA-KD surpasses other adaptive KDs, including logit-level adaptive LS [38] and sample-level adaptive approach PAD [46]. These results highlight EA-methods’ scalability on larger, more diverse datasets. Furthermore, Tab. 5 underscores EA-KD’s advantage in distilling valuable knowledge from vision transformer-based teachers to CNN students. Unlike standard KD, which plateaued when using DeiT-B as the teacher, EA-KD consistently improved student performance, mitigating the capacity gap and utilizing the guidance of stronger teachers more effectively.
@@
Table 7: LLM Distillation Results. Rouge-L [24] scores averaged over five seeds on each dataset are reported. SFT denotes a student supervised fine-tuned on the dataset.
@@ 4.3 Empirical Analysis
Ablation Study. Two experiment were performed to evaluate the sensitivity of $T^{\prime}$ and the effect of each reweighting components in $w_{\text{EA}}$. (i) Tab. 8 shows that a $T^{\prime}$ of 3 consistently delivers optimal performance across model combinations, highlighting its robustness in reflecting sample value. (ii) As shown in Tab. 9, $w_{\text{base}}$ enhances KD by emphasizing valuable samples under the teacher’s guidance. While $w_{\text{interact}}$ integrates student dynamics, it may introduce noise and dilute the teacher’s guidance, reducing effectiveness. However, when integrated into $w_{\text{EA}}$, $H^{\mathcal{S}}$ acts as a scaling factor for $H^{\mathcal{T}}$ (Eq. 12). This ensures early training, where $H^{\mathcal{S}}$ is mostly high and the scaling near 1, is mainly guided by $H^{\mathcal{T}}$. As training progresses, $H^{\mathcal{S}}$ adaptively adjusts and tailors the weighting, leading to superior performance.
@@
KD [11] 73.33 $\uparrow$ 75.14 $\uparrow$ $\uparrow$ 74.76 $\uparrow$ $\uparrow$ 75.46 $\uparrow$
@@
FCFD [26] 76.62 $\uparrow$ 77.50 $\uparrow$ $\uparrow$ 77.42 $\uparrow$ $\uparrow$ 77.44 $\uparrow$
Distillation Loss and Sample Value. In Fig. 5, we analyze the distribution of distillation loss across samples grouped by low- to high-entropy quartiles (Q1–Q4) for KD and EA-KD on CIFAR-100 and Tiny-ImageNet. Notably, KD loss is dominated by low-value samples throughout training on both datasets. The overshadowing effect in KD’s uniform distillation scheme hinders the transfer of critical knowledge from high-value samples, as discussed in Sec. 3, ultimately leading to more significant accuracy gaps (Fig. 2(a)). In contrast, EA-KD emphasizes the focus on valuable samples (Fig. 5), bringing holistic performance improvements across entropy segments (Fig. 2(a)).
Student’s Unique Perspective. Fig. 6 compares the t-SNE visualizations [41] of KD and EA-KD with the high $H^{\mathcal{S}}$ and $w_{\text{EA}}$ samples highlighted [^4] See Sec. B.1 for further t-SNE visualizations showing the student’s evolving sample focus over training epochs.. Similar to Fig. 2(b), high $H^{\mathcal{S}}$ samples also lie near decision boundaries in the KD-student. This suggests that, despite having a different view from the teacher (Fig. 3), the student’s evolving capacity enables $H^{\mathcal{S}}$ to capture the samples crucial for its own learning progress. In EA-KD, integrating both $H^{\mathcal{T}}$ and $H^{\mathcal{S}}$ in $w_{\text{EA}}$ captures both the teacher’s informed assessment and the student’s dynamic learning needs. As a result, EA-KD achieves superior class separability as reflected by a higher Calinski-Harabasz (CH) index.
Figure 5: Loss Distribution vs. Entropy Quartiles over Epochs. Low-value samples (Q1, Q2; blue) dominate the KD loss, whereas EA-KD places more focus on high-value ones (Q3, Q4; red).
Figure 6: t-SNE of Students. Top 10% high $H^{\mathcal{S}}$ samples for KD (left) and $w_{\text{EA}}$ samples for EA-KD (right) are highlighted with triangles. High $H^{\mathcal{S}}$ samples cluster near decision boundaries, underscoring their learning value for the student.
@@
The enhanced generalization in EA-DKD also extends to hyperparameter robustness. DKD was introduced to decouple target class KD (TCKD) and non-target class KD (NCKD) in vanilla KD, balancing them with hyperparameters $\alpha$ and $\beta$. While $\alpha$ remains stable around 1.0, adjusting $\beta$ from 1.0 to 10.0 leads to fluctuating performance [47], with a variance of 0.31 (see Fig. 8). Notably, EA-DKD effectively reduces this fluctuation to a variance of 0.10 and consistently improves the performance. This enhanced robustness may be attributed to the synergy between DKD and EA in handling different aspects of knowledge transfer: DKD prevents NCKD from being overshadowed by TCKD at the class level, while EA ensures valuable samples—often rich in NCKD—are not dominated by simpler ones at the sample level. Additionally, the dynamic focus in EA compensates for the static nature of $\beta$. Together, EA-DKD facilitates a nuanced and balanced knowledge transfer process, both class-wise and sample-wise.
Computational Efficiency. Fig. 1 illustrates the efficiency of EA-methods, showing significant performance improvements across various SOTA KD frameworks with negligible cost. This remarkable efficiency, combined with its streamlined integration into both logit- and feature-based KD methods, underscores the potential of our method as a versatile and practical enhancement for KD.
@@ 5 Conclusion
In this paper, we revisited existing KD methods from a novel perspective and revealed a key limitation in their inherent uniform distillation strategy, which often hinders the transfer of high-entropy samples that carry critical knowledge. To address this, we proposed EA-KD, a plug-and-play KD approach that dynamically reweights the distillation loss, directing the learning focus toward valuable samples. EA-KD consistently enhances representative KD baselines across image classification, object detection, and LLM distillation, all with negligible cost. We believe EA-KD showcases a great paradigm of the meticulous handling of knowledge transfer, adapting KD to the varying learning value of samples while accounting for the student’s evolving learning dynamics throughout the distillation process.
@@
Appendix A Additional Results and Comparison
A.1 Comparison with TTM and WTTM
TTM [48] removes student temperature in KD, revealing an inherent Rényi entropy regularizer that could improve generalization. WTTM further up-weights uncertain samples using the power sum of the teacher’s output, akin to our $w_{\text{base}}$ [^5] In fact, they also noted teacher entropy as a potential weighting factor, but left its systematic exploration for future work.. Notably, TTM’s Rényi entropy is an inherent effect of its structural modification, while EA-KD actively uses Shannon entropy to prioritize valuable samples. Moreover, unlike WTTM’s static teacher-based weighting, EA-KD dynamically adapts to the student’s learning (via $H^{\mathcal{S}}$), yielding stronger enhancement to TTM (Tab. A1). Finally, while their methods combine with feature-based KDs by adding and balancing their losses, EA-KD offers direct integration to both logit- and feature-based KDs through reweighting.
Table A1: TTM, WTTM, and EA-TTM on CIFAR-100 over 5 runs. Unlike WTTM, EA-TTM consistently improves TTM, highlighting the superiority of dynamic over static weighting.
Teacher Student TTM WTTM EA-TTM (Ours)
[…]
A.2 MS-COCO
Tab. A2 reports additional results for MS-COCO, where EA-DKD is compared against DKD [47] using a ResNet-101 to ResNet-18 distillation setting. The results show that EA-DKD consistently improves upon DKD across all evaluation metrics.
Table A2: More Results on MS-COCO.
[…]
DKD [39] 35.05 56.60 37.54
EA-DKD 35.16+0.11 56.75 +0.15 37.82 +0.28
A.3 LLM Distillation
Tab. A3 extends the comparison of EA-KD with Skewed KLD (SKLD) and SRKLD from the recent DistiLLM [15] method, under the same off-policy setting without pre-training corpus. EA-KD outperforms these stronger methods on most datasets, highlighting its overall effectiveness.
Table A3: More LLM Distillation Results.
Method Dolly SInst Vicuna S-NI UnNI Avg.
SKLD (DistiLLM) 24.03 10.66 14.70 17.99 21.18 17.71
SRKLD (DistiLLM) 24.48 10.35 14.88 16.53 19.68 17.19
EA-KD 24.95 10.59 16.41 18.27 21.46 18.34
Appendix B Additional Analysis and Visualizations
This section provides additional analysis to complement the main paper. It includes a detailed examination of high student entropy samples, loss landscape comparisons for KD and EA-KD, and t-SNE visualizations for various KD frameworks and their EA-enhanced variants.
B.1 Analysis of High Student Entropy Samples
Complementary to the teacher entropy analysis in Sec. 1, this section emphasizes the critical role of high student entropy samples in the adaptive reweighting process of EA-KD. As shown in Fig. 1(a), these samples correlate with larger teacher-student accuracy gaps similar to the teacher entropy segments (Fig. 2(a)). Additionally, Fig. 1(b) presents the students’ t-SNE visualizations across training epochs for KD and EA-KD, with high $H^{\mathcal{S}}$ and high $w_{\text{EA}}$ samples highlighted, respectively. Notably, these samples also cluster near decision boundaries similar to the teacher’s t-SNE (Fig. 2(b)). Unlike the teacher, however, the student’s top-entropy samples shift dynamically over training. For instance, in KD, the yellow class at epoch 40 is located near the center of most clusters and contains numerous high-entropy samples. By epoch 120, high-entropy samples are more prevalent in the green and purple classes as they shift closer to the center. With this dynamic integrated into $w_{\text{EA}}$, EA-KD continuously adapts to the student’s evolving learning needs throughout training, complementing the static nature of $w_{\text{base}}$ and fostering a tailored knowledge transfer.
@@ B.1 Analysis of High Student Entropy Samples
Figure B1: High Student Entropy Analysis in KD and EA-KD. (a) Higher student entropy samples correlate with larger accuracy gaps in KD. EA-KD demonstrates closer alignment with the teacher, particularly in high-entropy regions. (b) High $H^{\mathcal{S}}$ samples consistently cluster near decision boundaries, reflecting the student’s real-time learning needs. EA-KD adapts to this dynamic, achieving enhanced class separability over epochs.
@@
Figure B2: Loss Surface and Differences in Area@1.6 for KD and EA-KD Students Across Epochs. Contour plots illustrate the loss landscapes of KD (first row) and EA-KD (second row) across training epochs. The line plot (lower right) tracks the differences in Area@1.6 between the teacher and students over epochs. EA-KD exhibits a more stable and robust loss surface, with greater Area@1.6 earlier in training, signifying a more efficient learning process.
B.2 Loss Landscape Analysis for KD and EA-KD
We further analyze the robustness of KD and EA-KD by examining their loss landscapes across training epochs. As shown in Fig. B2, EA-KD achieves a consistently smoother loss surface compared to KD (e.g. $\sigma^{2}$ of 217.68 vs. 305.21 at epoch 240) and larger Area@1.6 values. Notably, KD’s Area@1.6 metric remains 0 at epochs 40 and 120, indicating the absence of low-loss regions. In contrast, EA-KD achieves values of 0.01 and 0.05, respectively, underscoring its improved learning efficiency by enabling the student to assimilate the key knowledge. These results highlight EA-KD’s ability to ensure a more efficient optimization, facilitating enhanced generalizability compared to standard KD.
B.3 t-SNE of KD Frameworks and EA-methods
The t-SNE visualizations of students from various KD frameworks and their EA-method variants, alongside the teacher’s representation, are shown in Fig. B3. EA-methods consistently exhibit more distinct and well-defined class clusters compared to their respective baselines, as evidenced by higher CH indices. Remarkably, the SOTA EA-MLD+LS achieves the highest CH index of 814.96, indicating its superior performance and closer alignment with the teacher. Furthermore, although EA-DKD ranks fourth in performance (Tab. 2), it demonstrates the second-best class separability, underscoring its enhanced robustness as discussed in Sec. 4.3. These findings highlight EA-KD’s versatility in enhancing class separability and improving knowledge transfer across diverse KD frameworks.